OpenStudy (anonymous):

All I need you to do is check if I did this essay right! Please, need help!!

OpenStudy (anonymous):

Taye and Mike are raking leaves to earn some money. Taye can rake 2 lawns that are about the same size in 3 hours. Mike can rake 2 lawns that are about the same in 4 hours. How long would it take both boys to work together to rake 2 lawns? Write and solve an equation for this situation. Explain how to set up the equation, using w = rt.

OpenStudy (anonymous):

Answer: rate of Taye= 2/3 lawns per hour rate of Mike=1/2 lawns per hour if they work together rate will be= 2/3+1/2=7/6 lawns per hour so needed time= 2/(7/6)= 12/7 hours= 1.714 hours